the width of a rectangle is 9x and the length is 3x + 11 solve for x then find the perimeter.

Answer by Earlsdon(6294)

P = 2(L+W) Substitute L = 3x+11 and W = 9x.
P = 2(3x+11+9x) Simplify.
P = 2(12x+11)
P = 24x+22
You can't find the perimeter (P) until you have the value of x.